# Performant Synchronization in Geo-Distributed Databases

### v1

@@ 6.5 Combining with other work
Combining with Compression Techniques. Compression is a widely used technique for reducing data volume. We select zlib [51, 22], one of the most popular compression libraries, to evaluate the performance of integrating compression with GeoCoCo. We conducted experiments on the YCSB workload, where the data transmission block size was set to 4 MB. Figure 16 reports the normalized makespan of one synchronization round across four configurations: Baseline, zlib, GeoCoCo, and GeoCoCo+zlib. Compression alone reduces the makespan by 54.13% compared to the baseline by shrinking data volume, while GeoCoCo achieves a larger reduction through WAN-aware path planning and semantic filtering. Combining zlib with GeoCoCo yields the lowest makespan, which is 33.62% of baseline. The results show that byte-level compression and GeoCoCo target complementary dimensions and can be effectively stacked.
@@ 7 RELATED WORK
Numerous research efforts have been dedicated to enhancing transmission performance in distributed database systems [44, 13, 37, 45, 57, 66, 64, 38, 24, 35, 36, 69, 86, 25, 47, 4, 8, 77, 76, 63, 60, 83, 85, 62, 32, 51, 22], primarily focusing on 1) cross transaction-transmission scheduling, 2) transmission optimization, 3) data transmission reduction.
@@
Cross Transaction-Transmission Scheduling. The goal of data transmission in a distributed database is to ensure the correctness of transactions. Considering the hardware feature of network devices, many researchers have dedicated for optimizing the scheduler between transaction and transmission [44, 13, 37, 45, 57, 66, 75, 64, 38, 85, 62, 32]. For example, AlNiCo [44] leverages FPGA-based smart NICs to schedule transaction requests, thereby reducing contention among multiple CPU cores. Hydra [13] uses programmable switches (P4) to perform transaction ordering at the network layer, offloading the ordering logic from servers to reduce their processing burden. SwitchTx [45] offloads distributed transaction coordination to a tree of programmable switches. It reorders in-flight transaction messages based on semantic priorities and tightly couples admission control with congestion control to optimize traffic. Xenic [57] serves as a smart-NIC-optimized transaction engine. It adopts an asynchronous aggregation execution model with custom NIC logic to maximize both network and core efficiency, thereby reducing commit latency. Compared to approaches that integrate transaction processing with network-level scheduling using custom hardware such as programmable switches or smart NICs, GeoCoCo does not rely on any dedicated network devices or modify the underlying transaction scheduling mechanism. This design allows GeoCoCo to be seamlessly integrated with existing concurrency control schemes at minimal integration cost, providing complementary benefits.
@@
Cross-region transmission optimization. Cross-region transmission optimization improves communication efficiency across geographically distributed data centers. Compared to intra-datacenter networks, WANs exhibit higher latency, fluctuating bandwidth, and dynamic link conditions, posing unique challenges for efficient data transfer. To address these issues, modern systems adopt optimizations across multiple layers [24, 35, 36, 69, 86, 25, 47, 4, 8]. At the application layer, RPC frameworks like gRPC [24, 35] leverage HTTP/2 to enable long-lived, pipelined streaming. At the transport layer, QUIC—built on UDP—integrates TLS [36, 69, 86], 0-RTT handshake [25], multiplexing, and connection migration, making it well-suited for WAN conditions. At the network level, protocols like TACK [47] reduce ACK overhead on high-latency paths, while Copa [4] balances latency and fairness in congestion control. GeoCoCo differs from these designs in two key ways: it integrates real-time latency into dynamic planning and optimizes transmission at a topology level, rather than treating flows in isolation. This enables GeoCoCo to adapt to live network conditions and optimize WAN-wide performance.
@@
Data transmission reduction. Reducing network traffic to improve transmission efficiency constitutes a classic research direction [77, 76, 63, 60, 83, 81, 52, 61, 82, 80, 10, 5, 51, 22, 84, 11, 12, 78]. To minimize data transfer, geo-replicated systems typically leverage a combination of compression [83], deduplication [77, 76], delta encoding [63], and aggregation [81, 52, 61, 82, 80, 10, 5]. In particular, delta encoding transmits only the differences or compressed change logs rather than full records. For instance, dbDedup [76] is an online deduplication scheme for databases that leverages similarity-aware, record-level delta encoding to reduce both storage usage and replication traffic. PHLIP [63], etc. introduces a wide-area data replication architecture that integrates stream-informed delta compression into existing deduplication systems, effectively addressing bandwidth limitations in remote backup scenarios. Cougar [81], etc. pushes operators to intermediate nodes for in-network aggregation, avoiding the transmission of raw readings from all sources and significantly reducing network traffic. GeoCoCo represents the first work applying these data reduction principles at the network layer of distributed databases. Its novel approach significantly reduces network load while strictly maintaining transaction correctness, improving transmission performance by 40.3% by reducing bandwidth usage.
@@ 8 CONCLUSION
This work was partly inspired by early-stage ideas that arose from discussions with Huawei experts Yang Ren and Jiaqi Liang, and was incubated with support from our joint collaboration project with Huawei. We also thank Zili Meng and Feng Zhang for their valuable guidance.
This work is supported by the National Natural Science Foundation of China (NSFC) under Projects No. 62202473, No. 62572473, No. 62272466, No. 62441230, No. 62322213, No. 62441230, No. 62461146205 and No. 62461146205, Shuimu Scholar Fellowship of Tsinghua University, as well as by the Postdoctoral Fellowship Program of CPSF under Grant No. GZC20251044.

### v3

@@ 6.5 Combining with other work
Combining with Compression Techniques. Compression is a widely used technique for reducing data volume. We select zlib [54, 23], one of the most popular compression libraries, to evaluate the performance of integrating compression with GeoCoCo. We conducted experiments on the YCSB workload, where the data transmission block size was set to 4 MB. Figure 16 reports the normalized makespan of one synchronization round across four configurations: Baseline, zlib, GeoCoCo, and GeoCoCo+zlib. Compression alone reduces the makespan by 54.13% compared to the baseline by shrinking data volume, while GeoCoCo achieves a larger reduction through WAN-aware path planning and semantic filtering. Combining zlib with GeoCoCo yields the lowest makespan, which is 33.62% of the baseline. The results show that byte-level compression and GeoCoCo target complementary dimensions and can be effectively stacked.
@@ 7 RELATED WORK
Numerous research efforts have been dedicated to enhancing transmission performance in distributed database systems [33, 91, 45, 14, 38, 46, 60, 69, 67, 39, 25, 36, 37, 72, 49, 89, 26, 50, 4, 9, 80, 79, 66, 63, 86, 88, 65, 33, 54, 23], primarily focusing on 1) cross transaction-transmission scheduling, 2) transmission optimization, 3) data transmission reduction.
@@
Cross Transaction-Transmission Scheduling. The goal of data transmission in a distributed database is to ensure the correctness of transactions. Considering the hardware feature of network devices, many researchers have dedicated for optimizing the scheduler between transaction and transmission [45, 14, 38, 46, 60, 69, 78, 67, 39, 88, 65, 33, 91]. For example, AlNiCo [45] leverages FPGA-based smart NICs to schedule transaction requests, thereby reducing contention among multiple CPU cores. Hydra [14] uses programmable switches (P4) to perform transaction ordering at the network layer, offloading the ordering logic from servers to reduce their processing burden. SwitchTx [46] offloads distributed transaction coordination to a tree of programmable switches. It reorders in-flight transaction messages based on semantic priorities and tightly couples admission control with congestion control to optimize traffic. Xenic [60] serves as a smart-NIC-optimized transaction engine. It adopts an asynchronous aggregation execution model with custom NIC logic to maximize both network and core efficiency, thereby reducing commit latency. Compared to approaches that integrate transaction processing with network-level scheduling using custom hardware such as programmable switches or smart NICs, GeoCoCo does not rely on any dedicated network devices or modify the underlying transaction scheduling mechanism. This design allows GeoCoCo to be seamlessly integrated with existing concurrency control schemes at minimal integration cost, providing complementary benefits.
@@
Cross-region transmission optimization. Cross-region transmission optimization improves communication efficiency across geographically distributed data centers. Compared to intra-datacenter networks, WANs exhibit higher latency, fluctuating bandwidth, and dynamic link conditions, posing unique challenges for efficient data transfer. To address these issues, modern systems adopt optimizations across multiple layers [25, 36, 37, 72, 89, 26, 50, 4, 9, 49]. At the application layer, RPC frameworks like gRPC [25, 36] leverage HTTP/2 to enable long-lived, pipelined streaming. At the transport layer, QUIC—built on UDP—integrates TLS [37, 72, 89], 0-RTT handshake [26], multiplexing, and connection migration, making it well-suited for WAN conditions. At the network level, protocols like TACK [50] reduce ACK overhead on high-latency paths, while Copa [4] balances latency and fairness in congestion control. GeoCoCo differs from these designs in two key ways: it integrates real-time latency into dynamic planning and optimizes transmission at a topology level, rather than treating flows in isolation. This enables GeoCoCo to adapt to live network conditions and optimize WAN-wide performance.
@@
Data transmission reduction. Reducing network traffic to improve transmission efficiency constitutes a classic research direction [80, 79, 66, 63, 86, 84, 55, 64, 85, 83, 11, 5, 54, 23, 87, 12, 13, 7, 81]. To minimize data transfer, geo-replicated systems typically leverage a combination of compression [86], deduplication [80, 79], delta encoding [66], and aggregation [84, 55, 64, 85, 83, 11, 5, 47]. In particular, delta encoding transmits only the differences or compressed change logs rather than full records. For instance, dbDedup [79] is an online deduplication scheme for databases that leverages similarity-aware, record-level delta encoding to reduce both storage usage and replication traffic. PHLIP [66], etc. introduces a wide-area data replication architecture that integrates stream-informed delta compression into existing deduplication systems, effectively addressing bandwidth limitations in remote backup scenarios. Cougar [84], etc. pushes operators to intermediate nodes for in-network aggregation, avoiding the transmission of raw readings from all sources and significantly reducing network traffic. GeoCoCo represents the first work applying these data reduction principles at the network layer of distributed databases. Its novel approach significantly reduces network load while strictly maintaining transaction correctness, improving transmission performance by 40.3% by reducing bandwidth usage.
@@ 8 CONCLUSION
This work was partly inspired by early-stage ideas that arose from discussions with Huawei experts Yang Ren and Jiaqi Liang, and was incubated with support from our joint collaboration project with Huawei. We also thank Yunpeng Chai, Zili Meng and Feng Zhang for their valuable guidance.
This work is supported by the National Natural Science Foundation of China (NSFC) under Projects No.62202473, No.62572473, No.62272466, No.62441230, No.62322213, No.62441230, No.62461146205, No. 62461146205 and No.62461146205, Shuimu Scholar Fellowship of Tsinghua University, as well as by the Postdoctoral Fellowship Program of CPSF under Grant No.GZC20251044.